OpenStudy (anonymous):

find the integral of (1+e^x)/(1-e^x)

OpenStudy (ash2326):

(1+e^x)(1-e^x)=1-e^2x now integrating we get x-e^2x/2+c

OpenStudy (anonymous):

can you pls show me the step by step procedure?

OpenStudy (anonymous):

i think you forgot the fraction bar between the two quantities..

OpenStudy (ash2326):

oh sorry

OpenStudy (anonymous):

please help me

OpenStudy (ash2326):

(1+e^x)/(1-e^x) \[1/(1-e^x)+e^x/(1-e^x)\] now integrate both terms one by one in first term we divide numerator and denominator by e^x we get \[e^-x/e^-x-1\] now if we take (e^-x-1) as t then -e^-xdx=dt so e^-xdx=-dt we are left for first term as -dt/t integrate it we -ln(t) t=e^-x-1 so integration of first we get -ln(e^-x-1)

OpenStudy (ash2326):

now second term is e^x/(1-e^x) if we take 1-e^x=t then e^xdx=-dt so dt/t on integrating gives -ln (t) so second term's integration gives -ln(1-e^x) so complete answer is -ln(e^-x-1)-ln(1-e^x)+c

OpenStudy (ash2326):

did you get it?

OpenStudy (anonymous):

yes, thanks..I'm a little confused about that problem a while ago, thanks for the help

OpenStudy (anonymous):

\[ \int \frac{1+e^x}{1-e^x}\] \[\int \frac{1}{1-e^x}+\frac{e^x}{1-e^x}\] \[\int \frac{e^{-x}}{e^{-x}-1}+\frac{e^x}{1-e^x}\] \[-ln(e^{-x}-1)-ln(1-e^{x})+C\]